OpenStudy (anonymous):

In how many permutations can 7 flags be arranged taken 4 at a time?

hero (hero):

7P4

OpenStudy (anonymous):

\[ 7\times 6 \times 5 \times 4 = \cdots \]

OpenStudy (lalaly):

smart^

OpenStudy (anonymous):

lol, not really. See the name: FOOOOL :P

OpenStudy (lalaly):

haha it has a different meaning ill tell u later:P